3. (4 pts) Find the shortest distance from the surface xy + 3x + z2 = 9 to the origin.
Answer #1

d² = x² + y² + 72 solution 3) The given Surface equation is ny + 3x + 7²=9 Now, the distance from the origin to given Point (and For shortest distance take partial derivative with resped to x and y ethal to zero. (13) pa (x240°49-ny-30) Ja 2x-y-3 2xFrom equation @ Put x=2y in equation o, 2 x 2y - Y-3=0 3y=3 or, y = 1 Then, x=2 So, the required shortest distance is d =V@)2